### v1

# Gradient-free stochastic optimization for additive models

Abstract
We address the problem of zero-order optimization from noisy observations for an objective function satisfying the Polyak-Łojasiewicz or the strong convexity condition. Additionally, we assume that the objective function has an additive structure and satisfies a higher-order smoothness property, characterized by the Hölder family of functions. The additive model for Hölder classes of functions is well-studied in the literature on nonparametric function estimation, where it is shown that such a model benefits from a substantial improvement of the estimation accuracy compared to the Hölder model without additive structure. We study this established framework in the context of gradient-free optimization. We propose a randomized gradient estimator that, when plugged into a gradient descent algorithm, allows one to achieve minimax optimal optimization error of the order $dT^{-(\beta-1)/\beta}$, where $d$ is the dimension of the problem, $T$ is the number of queries and $\beta\geq 2$ is the Hölder degree of smoothness. We conclude that, in contrast to nonparametric estimation problems, no substantial gain of accuracy can be achieved when using additive models in gradient-free optimization.
1 Introduction
Additive modeling is a popular approach to dimension reduction in nonparametric estimation problems [28, 12, 30]. It consists of considering that the unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ to be estimated from the data has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, where $x_{j}$’s are the coordinates of $\boldsymbol{x}\in\mathbb{R}^{d}$ and $f_{j}$’s are unknown functions of one variable. The main property proved in the literature on additive models in nonparametric regression can be summarized as follows. If each of the functions $f_{j}$ is $\beta$-Hölder (see Definition 1 below) then the minimax rate of estimation of $f$, pointwise or in $L_{2}$-norm, is of the order $n^{-\beta/(2\beta+1)}$, where $n$ is the number of observations [28]. This is in contrast with the problem of estimating $\beta$-Hölder functions on $\mathbb{R}^{d}$ without any additive structure, since for such functions the minimax rate is known to be $n^{-\beta/(2\beta+d)}$ [26, 27, 14, 13]. Thus, there is a substantial improvement in the rate of estimation when passing from general to additive models in nonparametric regression setting.
In the present paper, we show that such a dimension reduction property fails to hold in the context of gradient-free optimization. We consider additive modeling in the problem of minimizing an unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ when only sequential evaluations of values of $f$ are available, corrupted with noise. We assume that $f$ is either strongly convex or satisfies the Polyak-Łojasiewicz (PL) condition [22, 18] and admits an additive representation as described above, where the components $f_{j}$ are $\beta$-Hölder.
The setting that we consider belongs to the family of gradient-free (or zero-order) stochastic optimization problems, for which a rich literature is now available, see [17, 6, 19, 23, 15, 25, 10, 20, 4, 2, 5, 1, 8, 21] and the references therein. These papers did not assume any additive structure of $f$. It was proved in [23] that the minimax optimal rate of the optimization error, when $f$ is $\beta$-Hölder with $\beta\geq 2$ and satisfies the quadratic growth condition, is of the order $T^{-(\beta-1)/\beta}$ as function of the number of sequential queries $T$, to within an unspecified factor depending on the dimension $d$. Further developments were devoted to exploring the dependency of the minimax rate on $d$ assuming that $f$ is $\beta$-Hölder with $\beta\geq 2$ and is either strongly convex [15, 25, 4, 3, 1, 21] or satisfies the PL condition [1, 9]. In the PL case, unconstrained minimization was studied while the strongly convex case was analyzed both in constrained and unconstrained settings. A considerable progress was achieved though the complete solution valid for all $\beta\geq 2$ is not yet available. There exists a minimax lower bound for the class of $\beta$-Hölder and strongly convex functions, which scales as $dT^{-(\beta-1)/\beta}$ (proved in [25] for $\beta=2$ and Gaussian noise and in [2] for all $\beta\geq 2$ and more general noise; see also [1] for a yet more general lower bound). Moreover, the same rate is attained for $\beta=2$ under general conditions on the noise (no independence or zero-mean assumption), see [2]. Thus, for $\beta=2$ we know that the minimax rate is of the order $d\sqrt{T}$. For $\beta>2$, the literature provides different dependencies of the upper bounds on $d$ determined by the geometry of the $\beta$-Hölder condition. Thus, for the Hölder classes defined by pointwise Taylor approximation the best known upper bound is of the order $d^{2-1/\beta}T^{-(\beta-1)/\beta}$ when strongly convex functions are considered [3, 21]. On the other hand, for the Hölder classes defined by tensor-type conditions one can achieve the rate $d^{2-2/\beta}T^{-(\beta-1)/\beta}$ both for strongly convex functions and for PL functions [1]. Finally, the recent paper [31], assuming again strong convexity, deals with the class of functions that admit a Lipschitz Hessian. This represents a type of Hölder condition for $\beta=3$ and the paper [31] derives the upper rate $dT^{-2/3}$. We note that the lower bound of [2] with the rate $dT^{-(\beta-1)/\beta}$ is valid for all the above mentioned Hölder classes since this lower bound is obtained on additive functions that belong to all of them. Thus, the rate $dT^{-(\beta-1)/\beta}$ appears to be minimax optimal not only for $\beta=2$ but also for $\beta=3$ under a suitable definition of $3$-Hölder class of strongly convex functions.
The main result of the present paper is to establish an upper bound with the rate $dT^{-(\beta-1)/\beta}$ for the optimization error in the noisy gradient-free setting over the class of $\beta$-Hölder functions satisfying the additive model and either the PL condition or the strong convexity condition. Together with the lower bound proved in [2], it implies that $dT^{-(\beta-1)/\beta}$ is the minimax optimal rate of the optimization error in this setting for all $\beta\geq 2$. This conclusion is quite surprising as it goes against the intuition acquired from the classical results on nonparametric estimation mentioned above. Indeed, it means that, at least for $\beta\in\{2,3\}$, there is no improvement in the rate neither in $T$ nor in $d$ when passing from the general $\beta$-Hölder model to the additive $\beta$-Hölder model. If any, an improvement can be obtained only in a factor independent of $T$ and $d$. Such a property may be explained by the fact that the optimization setting is easier than nonparametric function estimation in the sense that it aims at estimating a specific functional of the unknown $f$ (namely, its minimizer) rather than $f$ as a whole object. We also refer to a another somewhat similar fact in the gradient-free stochastic optimization setting. Specifically, there is no dramatic difference between the complexity of minimizing a strongly convex function with Lipschitz gradient, which corresponds to the case $\beta=2$ discussed above with the minimax rate $d\sqrt{T}$, and the complexity of minimizing a convex function with no additional properties, for which one can construct an algorithm converging with the rate between $d^{1.5}\sqrt{T}$ and $d^{1.75}\sqrt{T}$ (up to a logarithmic factor) as recently shown in [7].
2 Problem setup
Let $\Theta$ be a closed convex subset of $\mathbb{R}^{d}$. We consider the problem of minimizing an unknown function $f:\mathbb{R}^{d}\to\mathbb{R}$ over the set $\Theta$ based on noisy evaluations of $f$ at query points that can be chosen sequentially depending on the past observations. Specifically, we assume that at round $t\in\{1,\dots,T\}$ we can observe two noisy evaluations of $f$ at points $\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}\in\mathbb{R}^{d}$, i.e.,
$y_{t}=f(\boldsymbol{z}_{t})+\xi_{t},\quad\quad y^{\prime}_{t}=f(\boldsymbol{z}^{\prime}_{t})+\xi^{\prime}_{t},$
where $\xi_{t},\xi^{\prime}_{t}$ are scalar noise variables and the query points $\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}$ can be chosen depending on $\{\boldsymbol{z}_{i},\boldsymbol{z}^{\prime}_{i},y_{i},y^{\prime}_{i}\}_{i=1}^{t-1}$ and on a suitable randomization.
Throughout the paper we assume that $f$ follows the additive model
$f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j}),$
where $x_{j}$’s are the coordinates of $\boldsymbol{x}\in\mathbb{R}^{d}$ and $f_{j}$’s are unknown functions of one variable.
We assume that each of the functions $f_{j}:\mathbb{R}\to\mathbb{R}$, $j=1,\dots,d$, belongs to the class of $\beta$-Hölder functions specified by the following definition, with $\beta\geq 2$.
Definition 1.
For $\beta>0$, $L>0$, denote by ${\cal F_{\beta}}(L)$ the set of all functions $f:\mathbb{R}\to\mathbb{R}$ that are $\ell=\lfloor\beta\rfloor$ times differentiable and satisfy, for all $x,z\in\mathbb{R}$, the condition
$\bigg{|}f(z)-\sum_{m=0}^{\ell}\frac{1}{m!}f^{(m)}(x)(z-x)^{m}\bigg{|}\leq L|z-x|^{\beta},$ (1)
where $f^{(m)}$ is the $m$-th derivative of $f$ and $\lfloor\beta\rfloor$ is the largest integer less than $\beta$. Elements of the class ${\cal F_{\beta}}(L)$ are referred to as $\beta$-Hölder functions.
If $\beta>2$ the fact that $f_{j}\in{\cal F_{\beta}}(L)$ does not imply that $f_{j}\in{\cal F}_{2}(L)$, however we will need the latter condition as well. It will be convenient to use it in a slightly different form given by the next definition.
Definition 2.
The function $f:\mathbb{R}\to\mathbb{R}$ is called ${\bar{L}}$-smooth if it is differentiable on $\mathbb{R}$ and there exists $\bar{L}>0$ such that, for every $x,x^{\prime}\in\mathbb{R}$, it holds that
$|f^{\prime}(x)-f^{\prime}(x^{\prime})|\leq{\bar{L}}|x-x^{\prime}|.$
The class of all ${\bar{L}}$-smooth functions will be denoted by ${\cal F}^{\prime}_{2}(\bar{L})$.
We also assume that $f$ is either an $\alpha$-strongly convex function or an $\alpha$-PL function as stated in the next two definitions.
Definition 3.
Let $\alpha>0$. The function $f:\mathbb{R}^{d}\to\mathbb{R}$ is called an $\alpha$-Polyak-Łojasiewicz function (shortly, $\alpha$-PL function) if $f$ is differentiable on $\mathbb{R}^{d}$ and
$2\alpha\left(f(\boldsymbol{x})-\min_{\boldsymbol{z}\in\mathbb{R}^{d}}f(\boldsymbol{z})\right)\leq\left\lVert\nabla f(\boldsymbol{x})\right\rVert^{2}\quad\quad\text{for all }\boldsymbol{x}\in\mathbb{R}^{d},$
[…]
Functions satisfying the PL condition are not necessarily convex. The PL condition is a useful tool in optimization problems since it leads to linear convergence of the gradient descent algorithm without convexity as shown by Polyak [22]. For more details and discussion on the PL condition see [16].
Definition 4.
Let $\alpha>0$. The function $f:\mathbb{R}^{d}\to\mathbb{R}$ is called $\alpha$-strongly convex if $f$ is differentiable on $\mathbb{R}^{d}$ and
$f(\boldsymbol{x})-f(\boldsymbol{x}^{\prime})\leq\langle\nabla f(\boldsymbol{x}),\boldsymbol{x}-\boldsymbol{x}^{\prime}\rangle-\frac{\alpha}{2}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert^{2}\quad\quad\text{for all }\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}.$
In order to minimize $f$, we apply a version of projected gradient descent with estimated gradient defined as follows. Let $\{\eta_{t}\}_{t=1}^{T}$ be a sequence of positive numbers and let $\{\boldsymbol{g}_{t}\}_{t=1}^{T}$ be a sequence of random vectors. Consider any fixed $\boldsymbol{x}_{1}\in\mathbb{R}^{d}$ and let the vectors $\boldsymbol{x}_{t}$ for $t=2,\dots,T$ be defined by the recursion
$\boldsymbol{x}_{t+1}=\text{\rm Proj}_{\Theta}\left(\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t}\right),$ (2)
where $\text{\rm Proj}_{\Theta}(\cdot)$ is the Euclidean projection over $\Theta$.
In this paper, the gradient estimator $\boldsymbol{g}_{t}=(g_{t,1},\dots,g_{t,d})\in\mathbb{R}^{d}$ at round $t\in\{1,\dots,T\}$ of the algorithm is defined as follows. For given $\beta\geq 2$ and $\ell=\lfloor\beta\rfloor$, let $K:[-1,1]\to\mathbb{R}$ be a function such that
[…]
At each round $t$ of the algorithm, we generate a random vector $\boldsymbol{r}_{t}=(r_{t,1},\dots,r_{t,d})\in\mathbb{R}^{d}$, where the components $r_{t,j}$ are independent and distributed uniformly on $[-1,1]$. For $h_{t}>0$, we draw two noisy evaluations
$y_{t}=f(\boldsymbol{x}_{t}+h_{t}\boldsymbol{r}_{t})+\xi_{t},\quad\quad y^{\prime}_{t}=f(\boldsymbol{x}_{t}-h_{t}\boldsymbol{r}_{t})+\xi^{\prime}_{t},$
and, for $j\in\{1,\dots,d\}$, we define
$g_{t,j}=\frac{1}{2h_{t}}(y_{t}-y^{\prime}_{t}){K(r_{t,j})}.$ (4)
We set $\boldsymbol{g}_{t}=(g_{t,1},\dots,g_{t,d})$.
We assume that the noise variables $\xi_{t},\xi^{\prime}_{t}$ and the randomizing variables $r_{t,j}$ satisfy the following.
Assumption 1.
There exists $\sigma^{2}>0$ such that for all $t\in\{1,\dots,T\}$ the following holds.
  • (i)
    The random variables $r_{t,j}\sim U[-1,1]$, $j=1,\dots,d$, are independent of $\boldsymbol{x}_{t}$ and conditionally independent of $\xi_{t},\xi^{\prime}_{t}$ given $\boldsymbol{x}_{t}$,
  • (ii)
    $\mathbb{E}[\xi_{t}^{2}]\leq\sigma^{2}$, $\mathbb{E}[(\xi^{\prime}_{t})^{2}]\leq\sigma^{2}$.
Assumption 1 (i) can be considered not as a restriction but as a part of the definition of the algorithm dealing with the choice of the randomizing variables $r_{t,j}$. Randomizations are naturally chosen independent of all other sources of randomness. For the proofs we need even a weaker property that we state here as an assumption in order to refer to it in what follows. Note also that Assumption 1 does not require the noises $\xi_{t},\xi^{\prime}_{t}$ to have zero mean. Moreover, they can be non-random and we do not assume independence between these noises on different rounds of the algorithm. The fact that convergence of gradient-free algorithms can be achieved under general conditions on the noise of similar type, not requiring independence and zero means, dates back to [11, 24].
3 Statement of the results
In this section, we provide upper bounds on the optimization error of the algorithm defined in Section 2. First, we assume that function $f$ represented by the additive model satisfies the $\alpha$-PL condition. Note that imposing this condition on the sum $f$ does not imply that all the components $f_{j}$ are PL functions. When dealing with PL functions we consider the problem of unconstrained minimization and we introduce the notation
$f^{*}=\min_{\boldsymbol{x}\in\mathbb{R}^{d}}f(\boldsymbol{x}).$
In what follows, for any positive integer $n$ we denote by $[n]$ the set of positive integers less than or equal to $n$.
Theorem 1.
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-PL function. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of algorithm (2) with $\Theta=\mathbb{R}^{d}$, where the gradient estimators $\boldsymbol{g}_{t}$ are defined by (4), and
$\eta_{t}$ $=\min\Big{(}\frac{4}{\alpha t},\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}\Big{)},$
$\vspace{4mm}h_{t}$ $=\Big{(}\frac{3\bar{L}}{\alpha}\cdot\frac{\kappa\sigma^{2}}{L^{2}\kappa_{\beta}^{2}}\Big{)}^{\frac{1}{2\beta}}\cdot\begin{cases}t^{-\frac{1}{2\beta}}&\text{if $\eta_{t}=\frac{4}{\alpha t}$},\vspace{3mm}\\ T^{-\frac{1}{2\beta}}&\text{if $\eta_{t}=\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}$}.\end{cases}$
Then,
$\mathbf{E}\left[f(\boldsymbol{x}_{T})-f^{*}\right]\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}(f(\boldsymbol{x}_{1})-f^{*})+\frac{d}{\alpha}\left(\texttt{A}_{1}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{2}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right),$
where $\texttt{A}_{1},\texttt{A}_{2}>0$ depend only on $\beta$ and $\sigma^{2}$.
Corollary 1.
Under the conditions of Theorem 1, if $T\geq(\bar{L}/\alpha)((\bar{L}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\alpha)^{\frac{\beta}{2}}/L)$, then
$\mathbf{E}\left[f(\boldsymbol{x}_{T})-f^{*}\right]\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}(f(\boldsymbol{x}_{1})-f^{*})+\texttt{A}\frac{d}{\alpha}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}},$
where $\texttt{A}>0$ depends only on $\beta$ and $\sigma^{2}$.
The bounds of Theorem 1 and Corollary 1 show how the rate of convergence depends on the parameters $T,d$ and $\alpha$ but also on the aspect ratio ${\bar{L}}/{\alpha}$. Moreover, in Corollary 1 the condition $T\geq Cd^{\beta/2}$, where $C>0$ is a constant, does not bring an additional restriction when $2\leq\beta\leq 3$ since the condition is weaker than giving the range of $T$, for which the bound makes sense. Indeed, for $2\leq\beta\leq 3$ the bound is greater than a constant independent of $T,d$ if $T\leq Cd^{\beta/2}$.
Next, we study the optimization error of the algorithm defined in Section 2 under the assumption that the objective function is $\alpha$-strongly convex. Unlike the case of PL functions, we consider here the problem of constrained minimization
$\min_{\boldsymbol{x}\in\Theta}f(\boldsymbol{x}),$
where $\Theta$ is a compact convex subset of $\mathbb{R}^{d}$.
Theorem 2.
Let $\alpha>0$, $\beta\geq 2$ and $\bar{L},L>0$. For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ has the form $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$, and that $f$ is an $\alpha$-strongly convex function such that $\max_{\boldsymbol{x}\in\Theta}\left\lVert\nabla f(\boldsymbol{x})\right\rVert\leq G$. Let Assumption 1 hold, and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of algorithm (2), where the gradient estimators $\boldsymbol{g}_{t}$ are defined by (4) with
$\eta_{t}=\frac{4}{\alpha(T+1)},\qquad h_{t}=\Big{(}\frac{3}{2t}\frac{\kappa\sigma^{2}}{\kappa_{\beta}^{2}L^{2}}\Big{)}^{\frac{1}{2\beta}}.$
Consider the weighted updates
$\bar{\boldsymbol{x}}_{T}=\frac{2}{T(T+1)}\sum_{t=1}^{T}t\boldsymbol{x}_{t}.$
Then, for any $\boldsymbol{x}\in\Theta$ we have
$\mathbf{E}\left[f(\bar{\boldsymbol{x}}_{T})-f(\boldsymbol{x})\right]\leq\frac{18G^{2}d}{\alpha T}+\frac{d}{\alpha}\left(\texttt{A}_{1}L^{\frac{2}{\beta}}+\texttt{A}_{2}d\bar{L}^{2}(LT)^{-\frac{2}{\beta}}\right)T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}_{1},\texttt{A}_{2}>0$ depend only on $\beta$ and $\sigma^{2}$.
Corollary 2.
Under the conditions of Theorem 2, if $T\geq(d\bar{L})^{\beta/2}L^{-2}$ then
$\mathbf{E}\Big{[}f(\bar{\boldsymbol{x}}_{T})-\min_{\boldsymbol{x}\in\Theta}f(\boldsymbol{x})\Big{]}\leq\texttt{A}\frac{d}{\alpha}T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}>0$ does not depend on $T,d$ and $\alpha$.
Similarly to Corollary 1, we may note that that for $2\leq\beta\leq 3$ the condition $T\geq Cd^{\beta/2}$ in Corollary 2, where $C>0$ is a constant, does not bring an additional restriction but indicates the meaningful range of $T$ since the bound is greater than a constant when $T\leq Cd^{\beta/2}$.
It follows from Corollary 2 and the proofs of the lower bounds in [2, 1] that, under non-restrictive conditions on the parameters of the problem, the rate $\frac{d}{\alpha}T^{-\frac{\beta-1}{\beta}}$ in Corollary 2 is minimax optimal on the class of additive functions $f$ satisfying the assumptions of Theorem 2. The lower bound that we need is not explicitly stated in [2, 1] but follows immediately from the proofs in those papers since the lower bounds in [2, 1] are obtained on additive functions. For completeness, we provide here the statement of the lower bound for additive functions based on [1].
Consider all strategies of choosing the query points as $\boldsymbol{z}_{t}=\Phi_{t}\big{(}\left(\boldsymbol{z}_{i},y_{i}\right)_{i=1}^{t-1},\left(\boldsymbol{z}^{\prime}_{i},y^{\prime}_{i}\right)_{i=1}^{t-1},\tau_{t}\big{)}$ and $\boldsymbol{z}^{\prime}_{t}=\Phi^{\prime}_{t}\big{(}\left(\boldsymbol{z}_{i},y_{i}\right)_{i=1}^{t-1},\left(\boldsymbol{z}^{\prime}_{i},y^{\prime}_{i}\right)_{i=1}^{t-1},\tau_{t}\big{)}$ for $t\geq 2$, where $\Phi_{t}$’s and $\Phi^{\prime}_{t}$’s are measurable functions, $\boldsymbol{z}_{1},\boldsymbol{z}^{\prime}_{1}\in\mathbb{R}^{d}$ are any random variables, and $\{\tau_{t}\}$ is a sequence of random variables with values in a measurable space $(\mathcal{Z},\mathcal{U})$, such that $\tau_{t}$ is independent of $\big{(}\left(\boldsymbol{z}_{i},y_{i}\right)_{i=1}^{t-1},\left(\boldsymbol{z}^{\prime}_{i},y^{\prime}_{i}\right)_{i=1}^{t-1}\big{)}$. We denote by $\Pi_{T}$ the set of all such strategies of choosing query points up to $t=T$. The class $\Pi_{T}$ includes the sequential strategy of the algorithm of Section 2 with the gradient estimator (4). In this case, $\tau_{t}=\boldsymbol{r}_{t}$, $\boldsymbol{z}_{t}=\boldsymbol{x}_{t}+h_{t}\boldsymbol{r}_{t}$ and $\boldsymbol{z}^{\prime}_{t}=\boldsymbol{x}_{t}-h_{t}\boldsymbol{r}_{t}$.
The lower bound of [1] that we are using here is proved under the following assumption on the noises $(\xi,\xi^{\prime}_{t})$. Let $H^{2}(\cdot,\cdot)$ be the squared Hellinger distance defined, for two probability measures $\mathbf{P},\mathbf{P}^{\prime}$ on a measurable space $(\Omega,\mathcal{A})$, as
$H^{2}(\mathbf{P},\mathbf{P}^{\prime})\triangleq\int(\sqrt{\,\mathrm{d}\mathbf{P}}-\sqrt{\,\mathrm{d}\mathbf{P}^{\prime}})^{2}\enspace.$
Assumption 2.
For every $t\geq 1$, the following holds:
  • The cumulative distribution function $F_{t}:\mathbb{R}^{2}\to\mathbb{R}$ of random variable $(\xi_{t},\xi^{\prime}_{t})$ is such that
    $H^{2}(P_{F_{t}(\cdot,\cdot)},P_{F_{t}(\cdot+v,\cdot+v)})\leq I_{0}v^{2}\,,\quad\quad|v|\leq v_{0}\enspace,$ (5)
    for some $0<I_{0}<\infty$, $0<v_{0}\leq\infty$. Here, $P_{F(\cdot,\cdot)}$ denotes the probability measure corresponding to the c.d.f. $F(\cdot,\cdot)$.
  • The random variable $(\xi_{t},\xi^{\prime}_{t})$ is independent of $((\boldsymbol{z}_{i},y_{i})_{i=1}^{t-1},(\boldsymbol{z}^{\prime}_{i},y^{\prime}_{i})_{i=1}^{t-1},\tau_{t})$.
Let $\Theta=\{\boldsymbol{x}\in\mathbb{R}^{d}\,:\,\left\lVert\boldsymbol{x}\right\rVert\leq 1\}$. Fix $\alpha,L,\bar{L}>0,G>\alpha$, $\beta\geq 2$, and denote by $\mathcal{F}$ the set of all functions $f$ that satisfy the assumptions of Theorem 2 and attain their minimum over $\mathbb{R}^{d}$ in $\Theta$.
Theorem 3.
Let $\Theta=\{\boldsymbol{x}\in\mathbb{R}^{d}\,:\,\left\lVert\boldsymbol{x}\right\rVert\leq 1\}$ and let Assumption 2 hold. Assume that $\alpha>T^{-1/2+1/\beta}$ and $T\geq d^{\beta}$. Then, for any estimator $\tilde{\boldsymbol{x}}_{T}$ based on the observations $((\boldsymbol{z}_{t},y_{t}),(\boldsymbol{z}^{\prime}_{t},y^{\prime}_{t}),t=1,\dots,T)$, where $((\boldsymbol{z}_{t},\boldsymbol{z}^{\prime}_{t}),t=1,\dots,T)$ are obtained by any strategy in the class $\Pi_{T}$ we have
$\sup_{f\in\mathcal{F}}\mathbf{E}\big{[}f(\tilde{\boldsymbol{x}}_{T})-\min_{\boldsymbol{x}\in\Theta}f(\boldsymbol{x})\big{]}\geq C\frac{d}{\alpha}T^{-\frac{\beta-1}{\beta}},$ (6)
where $C>0$ is a constant that does not depend of $T,d$, and $\alpha$.
Theorem 3 follows immediately from the proof of Theorem 22 in [1] since the family of functions used there belongs to the class $\mathcal{F}$. The lower bound of Theorem 22 in [1] has the form
$C\min\left(\max(\alpha,T^{-1/2+1/\beta}),\frac{d}{\sqrt{T}},\frac{d}{\alpha}T^{-\frac{\beta-1}{\beta}}\right),$
which reduces to $C\frac{d}{\alpha}T^{-\frac{\beta-1}{\beta}}$ under the assumptions on $T,d$ and $\alpha$ used in Theorem 3.
4 Proofs
We start by proving some auxiliary lemmas.
Lemma 1.
Let $f:\mathbb{R}^{d}\to\mathbb{R}$ be a differentiable function such that $\left\lVert\nabla f(\boldsymbol{x})-\nabla f(\boldsymbol{x}^{\prime})\right\rVert\leq\bar{L}\left\lVert\boldsymbol{x}-\boldsymbol{x}^{\prime}\right\rVert$ for all $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$, where $\bar{L}>0$. Let Assumption 1 hold and let $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ be the updates of algorithm (2) with $\Theta=\mathbb{R}^{d}$. Then, for all $t\in[T]$ we have
$\mathbf{E}\left[f(\boldsymbol{x}_{t+1})|\boldsymbol{x}_{t}\right]$ $\leq f(\boldsymbol{x}_{t})-\frac{\eta_{t}}{2}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}+\frac{\eta_{t}}{2}\left\lVert\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]-\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}+\frac{\bar{L}\eta_{t}^{2}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right].$ (7)
Proof.
The assumption on $f$ implies that
$f(\boldsymbol{x}_{t+1})$ $=f(\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t})\leq f(\boldsymbol{x}_{t})-\eta_{t}\langle\nabla f(\boldsymbol{x}_{t}),\boldsymbol{g}_{t}\rangle+\frac{\bar{L}\eta_{t}^{2}}{2}\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}.$
By adding and subtracting $\eta_{t}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}$ we obtain
$f(\boldsymbol{x}_{t+1})$ $\leq f(\boldsymbol{x}_{t})-\eta_{t}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}-\eta_{t}\langle\nabla f(\boldsymbol{x}_{t}),\ \boldsymbol{g}_{t}-\nabla f(\boldsymbol{x}_{t})\rangle+\frac{\bar{L}\eta_{t}^{2}}{2}\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}.$
Taking the conditional expectation gives
$\mathbf{E}\left[f(\boldsymbol{x}_{t+1})|\boldsymbol{x}_{t}\right]$ $\leq f(\boldsymbol{x}_{t})-\eta_{t}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}-\eta_{t}\langle\nabla f(\boldsymbol{x}_{t}),\ \mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]-\nabla f(\boldsymbol{x}_{t})\rangle+\frac{\bar{L}\eta_{t}^{2}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]$
$\leq f(\boldsymbol{x}_{t})-\eta_{t}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}+\eta_{t}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert\left\lVert\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]-\nabla f(\boldsymbol{x}_{t})\right\rVert+\frac{\bar{L}\eta_{t}^{2}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right].$
The lemma follows by using the inequality $2ab\leq a^{2}+b^{2}$, $\forall a,b\in\mathbb{R}$. ∎
Lemma 2.
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{\beta}(L)$, where $\beta\geq 2$ and $L>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1(i) hold. Then, for all $t\in[T]$ we have
$\left\lVert\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]-\nabla f(\boldsymbol{x}_{t})\right\rVert\leq\kappa_{\beta}L\sqrt{d}h_{t}^{\beta-1}.$
Proof.
Using Assumption 1(i) for any $j\in[d]$ and $t\in[T]$ we have $\mathbf{E}(K(r_{t,j}))=0$, $\mathbf{E}\left[\xi_{t}K(r_{t,j})|\boldsymbol{x}_{t}\right]=0$ and $\mathbf{E}\left[\xi^{\prime}_{t}K(r_{t,j})|\boldsymbol{x}_{t}\right]=0$. Thus,
$\mathbf{E}\left[g_{t,j}|\boldsymbol{x}_{t}\right]$ $=\frac{1}{2h_{t}}\mathbf{E}\left[(f_{j}(x_{t,j}+h_{t}r_{t,j})-f_{j}(x_{t,j}-h_{t}r_{t,j}))K(r_{t,j})|\boldsymbol{x}_{t}\right]$
$\qquad+\frac{1}{2h_{t}}\sum_{m\neq j}\mathbf{E}\left[(f_{m}(x_{t,m}+h_{t}r_{t,m})-f_{m}(x_{t,m}-h_{t}r_{t,m}))K(r_{t,j})|\boldsymbol{x}_{t}\right]$
$=\frac{1}{2h_{t}}\mathbf{E}\left[(f_{j}(x_{t,j}+h_{t}r_{t,j})-f_{j}(x_{t,j}-h_{t}r_{t,j}))K(r_{t,j})|\boldsymbol{x}_{t}\right],$
where we have used the fact that $\mathbf{E}(K(r_{t,j}))=0$ and $r_{t,j}$ is independent of $r_{t,m},$ for $m\neq j$, and of $\boldsymbol{x}_{t}$. From Taylor expansion we obtain that
$\frac{1}{2h_{t}}\left(f_{j}(x_{t,j}+h_{t}r_{t,j})-f_{j}(x_{t,j}-h_{t}r_{t,j})\right)$ $=f^{\prime}_{j}(x_{t,j})r_{t,j}+\frac{1}{h_{t}}\sum_{1\leq m\leq\ell,\,m\text{ odd}}\frac{h_{t}^{m}}{m!}f^{(m)}_{j}(x_{t,j})r_{t,j}^{m}$
$\quad\quad+\frac{R(h_{t}r_{t,j})-R(-h_{t}r_{t,j})}{2h_{t}},$
where $|R(-h_{t}r_{t,j})|,|R(h_{t}r_{t,j})|\leq L|r_{t,j}|^{\beta}h_{t}^{\beta}$. Multiplying both sides of this inequality by $K(r_{t,j})$ and taking the conditional expectation implies
$|\mathbf{E}\left[g_{t,j}|x_{t,j}\right]-f^{\prime}_{j}(x_{t,j})|\leq L\mathbf{E}\left[|r_{t,j}|^{\beta}K(r_{t,j})\right]h_{t}^{\beta-1}=\kappa_{\beta}Lh_{t}^{\beta-1},$
and thus the result of the lemma. ∎
Lemma 3.
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})$, where $\bar{L}>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1 hold. Then, for all $t\in[T]$ we have
$\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]$ $\leq\frac{3}{2}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right).$
Proof.
For $i\in[d]$ we define
$G_{i}=f_{i}(x_{t,i}+h_{t}r_{t,i})-f_{i}(x_{t,i}-h_{t}r_{t,i}).$
We have
$\mathbf{E}\left[g^{2}_{t,j}|\boldsymbol{x}_{t}\right]$ $=\frac{1}{4h_{t}^{2}}\mathbf{E}\left[\left(\sum_{i=1}^{d}G_{i}+\xi_{t}-\xi^{\prime}_{t}\right)^{2}K^{2}(r_{t,j})|\boldsymbol{x}_{t}\right]$
Note that $r_{t,i}$ and $-r_{t,i}$ have the same distribution. Therefore, $\mathbf{E}[G_{i}|\boldsymbol{x}_{t}]=0$ and we can write
$\mathbf{E}\left[g^{2}_{t,j}|\boldsymbol{x}_{t}\right]$ $\leq\frac{3}{4h_{t}^{2}}\mathbf{E}\left[\left(\left(\sum_{i=1}^{d}G_{i}\right)^{2}+\xi_{t}^{2}+(\xi^{\prime}_{t})^{2})\right)K^{2}(r_{t,j})|\boldsymbol{x}_{t}\right]$
$\leq\frac{3}{4h_{t}^{2}}\mathbf{E}\left[\sum_{i=1}^{d}G_{i}^{2}K^{2}(r_{t,j})+\sum_{i,k=1,i\neq k}^{d}G_{i}G_{k}K^{2}(r_{t,j})|\boldsymbol{x}_{t}\right]+\frac{3\sigma^{2}\kappa}{2h_{t}^{2}}$
$=\frac{3}{4h_{t}^{2}}\mathbf{E}\left[\sum_{i=1}^{d}G_{i}^{2}K^{2}(r_{t,j})|\boldsymbol{x}_{t}\right]+\frac{3\sigma^{2}\kappa}{2h_{t}^{2}}.$
Since $f_{i}\in\mathcal{F}_{2}^{\prime}(\bar{L})$ we have that, for all $i\in[d]$,
$G_{i}^{2}$ $=\left(\left(f_{i}(x_{t,i}+h_{t}r_{t,i})-f(x_{t,i})-f^{\prime}_{i}(x_{t,i})h_{t}r_{t,i}\right)-\left(f_{i}(x_{t,i}-h_{t}r_{t,i})-f(x_{t,i})+f^{\prime}_{i}(x_{t,i})h_{t}r_{t,i}\right)\right.$
+ $\qquad\left.+\,2f^{\prime}_{i}(x_{t,i})h_{t}r_{t,i}\right)^{2}$
$\leq 3\left(\left(f_{i}(x_{t,i}+h_{t}r_{t,i})-f(x_{t,i})-f^{\prime}_{i}(x_{t,i})h_{t}r_{t,i}\right)^{2}+\left(f_{i}(x_{t,i}-h_{t}r_{t,i})-f(x_{t,i})+f^{\prime}_{i}(x_{t,i})h_{t}r_{t,i}\right)^{2}\right.$
+ $\qquad\left.+\,4\left(f^{\prime}_{i}(x_{t,i})h_{t}r_{t,i}\right)^{2}\right)$
$\leq 3\left(\frac{\bar{L}^{2}}{2}h_{t}^{4}+4\left(f^{\prime}_{i}(x_{t,i})\right)^{2}h^{2}_{t}\right).$
Thus,
$\mathbf{E}\left[g^{2}_{t,j}|\boldsymbol{x}_{t}\right]$ $\leq\frac{9}{4}\kappa\left(\frac{d\bar{L}^{2}}{2}h_{t}^{2}+4\sum_{i=1}^{d}(f^{\prime}_{i}(x_{t,i}))^{2}\right)+\frac{3\sigma^{2}\kappa}{2h_{t}^{2}},$
which implies the lemma. ∎
Lemma 4.
For $j\in[d]$, let $f_{j}:\mathbb{R}\to\mathbb{R}$ be such that $f_{j}\in\mathcal{F}_{2}^{\prime}(\bar{L})\cap\mathcal{F}_{\beta}(L)$, where $\beta\geq 2$ and $\bar{L},L>0$. Assume that $f:\mathbb{R}^{d}\to\mathbb{R}$ satisfies the additive model $f(\boldsymbol{x})=\sum_{j=1}^{d}f_{j}(x_{j})$. Let Assumption 1 hold. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are updates of algorithm (2) with $\Theta=\mathbb{R}^{d}$ and with the gradient estimators defined by (4). Then, for all $t\in[T]$ we have
$\mathbf{E}\left[f(\boldsymbol{x}_{t+1})|\boldsymbol{x}_{t}\right]\leq f(\boldsymbol{x}_{t})-\frac{\eta_{t}}{2}\left(1-9\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa\eta_{t}\right)\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}+\frac{\eta_{t}}{2}d\left(L\kappa_{\beta}h_{t}^{\beta-1}\right)^{2}+\frac{3\bar{L}\eta_{t}^{2}}{4}\kappa d\left(\frac{\sigma^{2}}{h_{t}^{2}}+\frac{3\bar{L}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{4}h_{t}^{2}\right).$
Proof.
The result follows by combining Lemmas 1, 2 and 3. ∎
Lemma 5.
For $\alpha>0$ assume that $f$ is an $\alpha$-strongly convex function. Assume that $\{\boldsymbol{x}_{t}\}_{t=1}^{T}$ are the updates of algorithm (2). Then, for all $t\in[T]$ and $\boldsymbol{x}\in\Theta$, we have
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})$ $\leq\left(2\eta_{t}\right)^{-1}\left(\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}-\mathbf{E}\left[\left\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rVert^{2}|\boldsymbol{x}_{t}\right]\right)+\frac{1}{\alpha}\left\lVert\nabla f(\boldsymbol{x}_{t})-\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]\right\rVert^{2}$
$\quad\quad+\frac{\eta_{t}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]-\frac{\alpha}{4}\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}.$
Proof.
Fix $\boldsymbol{x}\in\Theta$. Then, by the contraction property of the Euclidean projection we have $\left\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rVert^{2}\leq\left\lVert\boldsymbol{x}_{t}-\eta_{t}\boldsymbol{g}_{t}-\boldsymbol{x}\right\rVert^{2}$. Equivalently,
$\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\rangle\leq(2\eta_{t})^{-1}(\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}-\left\lVert\boldsymbol{x}_{t+1}-\boldsymbol{x}\right\rVert^{2})+\frac{\eta_{t}}{2}\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}.$
Let $a_{t}=\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}$. Since $f$ is an $\alpha$-strongly convex function we have
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})$ $\leq\langle\nabla f(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}\rangle-\frac{\alpha}{2}a_{t}$
$=\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\rangle+\langle\nabla f(\boldsymbol{x}_{t})-\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\rangle-\frac{\alpha}{2}a_{t}$
$\leq(2\eta_{t})^{-1}(a_{t}-a_{t+1})+\langle\nabla f(\boldsymbol{x}_{t})-\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}\rangle+\frac{\eta_{t}}{2}\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}-\frac{\alpha}{2}a_{t}.$
Taking the conditional expectation given $\boldsymbol{x}_{t}$ and using the inequality $ab\leq a^{2}/\lambda+\lambda b^{2}/4$ valid for all $a,b\in\mathbb{R}$ and $\lambda>0$ we deduce that
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})$ $\leq(2\eta_{t})^{-1}(a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right])+\langle\nabla f(\boldsymbol{x}_{t})-\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right],\boldsymbol{x}_{t}-\boldsymbol{x}\rangle+\frac{\eta_{t}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]-\frac{\alpha}{2}a_{t}$
$\leq(2\eta_{t})^{-1}(a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right])+\left\lVert\nabla f(\boldsymbol{x}_{t})-\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]\right\rVert\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert+\frac{\eta_{t}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]-\frac{\alpha}{2}a_{t}$
$\leq(2\eta_{t})^{-1}(a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right])+\frac{1}{\alpha}\left\lVert\nabla f(\boldsymbol{x}_{t})-\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]\right\rVert^{2}+\frac{\eta_{t}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]-\frac{\alpha}{4}a_{t}.$
Proof of Theorem 1. Since $\eta_{t}\leq 1/(18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa)$ from Lemma 4 we have
$\mathbf{E}\left[f(\boldsymbol{x}_{t+1})|\boldsymbol{x}_{t}\right]$ $\leq f(\boldsymbol{x}_{t})-\frac{\eta_{t}}{4}\left\lVert\nabla f(\boldsymbol{x}_{t})\right\rVert^{2}+\frac{\eta_{t}}{2}d\left(L\kappa_{\beta}h_{t}^{\beta-1}\right)^{2}+\frac{3\bar{L}\eta_{t}^{2}}{4}\kappa d\left(\frac{\sigma^{2}}{h_{t}^{2}}+\frac{3\bar{L}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{4}h_{t}^{2}\right).$
Taking the expectation from both sides of this inequality and using the fact that $f$ is an $\alpha$-PL function we obtain
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{\eta_{t}\alpha}{2}\right)+\frac{\eta_{t}}{2}d\left(L\kappa_{\beta}h_{t}^{\beta-1}\right)^{2}+\frac{3\bar{L}\eta_{t}^{2}}{4}\kappa d\left(\frac{\sigma^{2}}{h_{t}^{2}}+\frac{3\bar{L}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{4}h_{t}^{2}\right),$ (8)
where $\delta_{t}=\mathbf{E}\left[f(\boldsymbol{x}_{t})-f^{*}\right]$. Let $T_{0}=\left\lfloor 72\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa/\alpha\right\rfloor$. Note that
$\eta_{t}=\begin{cases}\frac{1}{18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}\quad\quad&\text{if $t<T_{0}$},\vspace{3mm}\\ \frac{4}{\alpha t}\quad\quad&\text{if $t\geq T_{0}$}.\end{cases}$
We consider separately the cases $T>T_{0}$ and $T\leq T_{0}$.
If $T>T_{0}$, then for $t\geq T_{0}$ we have $\eta_{t}=4/(\alpha t)$ and we can write
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{2}{t}\right)+\frac{d}{\alpha t}\left(2\left(L\kappa_{\beta}h_{t}^{\beta-1}\right)^{2}+\frac{3\bar{L}}{\alpha t}\kappa\left(\frac{\sigma^{2}}{h_{t}^{2}}+\frac{3\bar{L}^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{4}h_{t}^{2}\right)\right).$
Substituting here $h_{t}=(\frac{3\bar{L}}{\alpha t}\cdot\frac{\kappa\sigma^{2}}{2L^{2}\kappa_{\beta}^{2}})^{\frac{1}{2\beta}}$ gives
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{2}{t}\right)+\frac{d}{\alpha t}\left(\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha t}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{t}\left(\frac{\bar{L}}{\alpha tL^{2}}\right)^{\frac{1}{\beta}}\right).$
where $\texttt{A}_{3}=6\,\left((2\kappa_{\beta}^{2}/3)(\kappa\sigma^{2})^{\beta-1}\right)^{1/\beta}$ and $\texttt{A}_{4}=9\left(3\kappa\sigma^{2}/(2\kappa_{\beta}^{2})\right)^{\frac{1}{\beta}}/4$. Since $T\geq T_{0}$, applying [1, Lemma 32] leads to the bound
$\delta_{T}$ $\leq\frac{2T_{0}}{T}\delta_{T_{0}+1}+\frac{d}{\alpha}\left(\frac{\beta\texttt{A}_{1}}{\beta+1}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\frac{\beta\texttt{A}_{2}}{2\beta+1}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
If $T_{0}=0$ then the proof is complete for the case $T>T_{0}$. Otherwise, for any $t\leq T_{0}$, we have $\eta_{t}=1/(18\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa)$ and $h_{t}=(\frac{3\bar{L}}{\alpha T}\cdot\frac{\kappa\sigma^{2}}{2L^{2}\kappa_{\beta}^{2}})^{\frac{1}{2\beta}}$. Note that
$\frac{4}{(T_{0}+1)\alpha}\leq\eta_{t}\leq\frac{4}{T_{0}\alpha},$
and from (8) we deduce that, for all $t\leq T_{0}$,
$\delta_{t+1}\leq\delta_{t}\left(1-\frac{2}{T_{0}+1}\right)+\frac{2d}{\alpha T_{0}}\left(\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha}\right)^{\frac{\beta-1}{\beta}}\left(T^{-\frac{\beta-1}{\beta}}+\frac{T^{\frac{1}{\beta}}}{T_{0}}\right)+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T_{0}}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Note that $1-2/(T_{0}+1)\leq 1$. Thus, by taking into account the definition of $T_{0}$ and the fact that $T>T_{0}$ we get
$\frac{2T_{0}}{T}\delta_{t+1}\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}\delta_{1}+\frac{4d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Therefore,
$\delta_{T}\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha T}\delta_{1}+\frac{d}{\alpha}\left(\texttt{A}_{1}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{2}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right),$
where $\texttt{A}_{1}=\texttt{A}_{3}\left(\beta/(\beta+1)+8\right)$ and $\texttt{A}_{2}=\texttt{A}_{4}\left(\beta/(2\beta+1)+4\right)$.
Consider now the case $T\leq T_{0}$. Then from (8) we get that, for all $t\leq T$,
$\delta_{t+1}$ $\leq\delta_{t}\left(1-\frac{2}{T_{0}+1}\right)+\frac{2d}{\alpha T_{0}}\left(\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha}\right)^{\frac{\beta-1}{\beta}}\left(T^{-\frac{\beta-1}{\beta}}+\frac{T^{\frac{1}{\beta}}}{T_{0}}\right)+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T_{0}}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Thus,
$\delta_{T+1}$ $\leq\delta_{1}\left(1-\frac{2}{T_{0}+1}\right)^{T}+\frac{2d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
Since $(1-\lambda)^{T}\leq\exp(-\lambda T)\leq 1/(\lambda T)$ for all $T,\lambda>0$, we deduce from the previous display that
$\delta_{T+1}$ $\leq\frac{T_{0}+1}{2T}\delta_{1}+\frac{2d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{\bar{L}}{\alpha T}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T}\left(\frac{\bar{L}}{\alpha TL^{2}}\right)^{\frac{1}{\beta}}\right).$
By using the fact that $\lambda+1\leq 2\lambda$ for all $\lambda\geq 1$ we finally get
$\delta_{T+1}$ $\leq\frac{2T_{0}}{T+1}\delta_{1}+\frac{2d}{\alpha}\left(2\texttt{A}_{3}\left(\frac{2\bar{L}}{\alpha(T+1)}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{4}\frac{2\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T+1}\left(\frac{2\bar{L}}{\alpha(T+1)L^{2}}\right)^{\frac{1}{\beta}}\right)$
$\leq\frac{144\bar{L}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}\kappa}{\alpha(T+1)}\delta_{1}+\frac{d}{\alpha}\left(\texttt{A}_{1}\left(\frac{\bar{L}}{\alpha(T+1)}\right)^{\frac{\beta-1}{\beta}}+\texttt{A}_{2}\frac{\bar{L}^{3}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}d}}{T+1}\left(\frac{\bar{L}}{\alpha(T+1)L^{2}}\right)^{\frac{1}{\beta}}\right),$
where $\texttt{A}_{1}=2^{\frac{3\beta-1}{\beta}}\texttt{A}_{3}$ and $\texttt{A}_{2}=2^{\frac{2\beta+1}{\beta}}\texttt{A}_{4}$.
Proof of Theorem 2. Fix $\boldsymbol{x}\in\Theta$. By Lemma 5 we have
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\leq\frac{a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right]}{2\eta_{t}}+\frac{1}{\alpha}\left\lVert\nabla f(\boldsymbol{x}_{t})-\mathbf{E}\left[\boldsymbol{g}_{t}|\boldsymbol{x}_{t}\right]\right\rVert^{2}+\frac{\eta_{t}}{2}\mathbf{E}\left[\left\lVert\boldsymbol{g}_{t}\right\rVert^{2}|\boldsymbol{x}_{t}\right]-\frac{\alpha}{4}a_{t},$
where $a_{t}=\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}$. Using Lemmas 2 and 3 and the assumption that $\max_{\boldsymbol{x}\in\Theta}\left\lVert\nabla f(\boldsymbol{x})\right\rVert\leq G$ we obtain
$f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\leq\frac{a_{t}-\mathbf{E}\left[a_{t+1}|\boldsymbol{x}_{t}\right]}{2\eta_{t}}+\frac{d}{\alpha}(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3\eta_{t}}{4}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right)-\frac{\alpha}{4}a_{t}.$
Let $b_{t}=\mathbf{E}\left[\left\lVert\boldsymbol{x}_{t}-\boldsymbol{x}\right\rVert^{2}\right]$. Using the definition $\eta_{t}=4/(\alpha(t+1))$ and taking the expectation we obtain
$\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{\alpha}{4}\left(t\left(b_{t}-b_{t+1}\right)-b_{t}\right)+\frac{d}{\alpha}(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3}{2\alpha(t+1)}\kappa d\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right).$
Summing up both sides of this inequality form $1$ to $T$ and using the fact that
$\sum_{t=1}^{T}t\left((t+1)\left(b_{t}-b_{t+1}\right)-b_{t}\right)\leq 0$
we find
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{d}{\alpha}\sum_{t=1}^{T}\left(t(\kappa_{\beta}Lh_{t}^{\beta-1})^{2}+\frac{3t}{2(t+1)}\kappa\left(\frac{3}{4}\left(d\bar{L}^{2}h_{t}^{2}+8G^{2}\right)+\frac{\sigma^{2}}{h_{t}^{2}}\right)\right).$
Substituting here $h_{t}=(\frac{3}{2t}\frac{\kappa\sigma^{2}}{\kappa_{\beta}^{2}L^{2}})^{\frac{1}{2\beta}}$ yields
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{9G^{2}dT}{\alpha}+\frac{d}{\alpha}\sum_{t=1}^{T}\left(\texttt{A}_{3}(L^{2}t)^{\frac{1}{\beta}}+\texttt{A}_{4}d\bar{L}^{2}\left(L^{2}t\right)^{-\frac{1}{\beta}}\right),$
where $\texttt{A}_{3}=2(3\kappa\sigma^{2}/2)^{\frac{\beta-1}{\beta}}{\kappa}_{\beta}^{\frac{2}{\beta}}$ and $\texttt{A}_{4}=(1/2)\left(3/2\right)^{\frac{2\beta+1}{\beta}}\kappa^{\frac{\beta+1}{\beta}}\left(\sigma/\kappa_{\beta}\right)^{\frac{2}{\beta}}$. Using the inequality $\sum_{t=1}^{T}t^{-\frac{1}{\beta}}\leq(\beta/(\beta-1))T^{\frac{\beta-1}{\beta}}$ we find
$\sum_{t=1}^{T}t\mathbf{E}\left[f(\boldsymbol{x}_{t})-f(\boldsymbol{x})\right]\leq\frac{9G^{2}dT}{\alpha}+\frac{d}{\alpha}\left(\texttt{A}_{3}L^{\frac{2}{\beta}}+\frac{\beta\texttt{A}_{4}}{\beta-1}d\bar{L}^{2}(LT)^{-\frac{2}{\beta}}\right)T^{\frac{\beta+1}{\beta}}.$
Dividing both sides of this inequality by $T(T+1)/2$ and applying Jensen’s inequality we finally get
$\mathbf{E}\left[f(\bar{\boldsymbol{x}}_{T})-f(\boldsymbol{x})\right]\leq\frac{18G^{2}d}{\alpha T}+\frac{d}{\alpha}\left(\texttt{A}_{1}L^{\frac{2}{\beta}}+\texttt{A}_{2}d\bar{L}^{2}(LT)^{-\frac{2}{\beta}}\right)T^{-\frac{\beta-1}{\beta}},$
where $\texttt{A}_{1}=2\texttt{A}_{3}$ and $\texttt{A}_{2}=2\beta\texttt{A}_{4}/(\beta-1)$.
Disclosure of Funding
The research of Arya Akhavan was funded by UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number EP/Y028333/1].